10.35 contains 0.119 mol of gas and has a volume of 2.83 L ½ 2.78
Answer #1

Part A

after addition of 0.117 mole of gas final mole = 0.119 + 0.117 = 0.236 mole

we know avogadro's law volume of gas directly proportional to no. of mole of gas at constant tempreture and pressure.

V \alpha n

another form of this equation

V1/n1 = V2/n2

where V1 = initial volume = 2.83 L

V2 = final volume = ?

n1 = initial mole = 0.119 mole

n2 = final mole = 0.236 mole

We can wright above eqation as

V2 = V1n2/n1

Substitute value

V2 = 2.83 X 0.236 / 0.119 = 5.61 L

final volume of gas = 5.61 L